OpenStudy (anonymous):

If "a" be a non zero complex number and z be any complex number then define a^z. Also what is the principal value of a^z If "a" be a non zero complex number and z be any complex number then define a^z. Also what is the principal value of a^z @Mathematics

OpenStudy (perl):

take the log of of this

OpenStudy (anonymous):

I know how to take the log and then solve it. But I don't understand, what they mean by defintion

OpenStudy (perl):

what are in your notes

OpenStudy (anonymous):

Its a question in our examination from past years

OpenStudy (anonymous):

and it says "If "a" be a non zero complex number and z be any complex number then define a^z."

OpenStudy (anonymous):

I really don't undersand how to define this..

OpenStudy (anonymous):

What do you think?

OpenStudy (anonymous):

But these are complex number..

OpenStudy (perl):

yes, it says complex

OpenStudy (perl):

For complex numbers a and b with a ≠ 0, the notation ab is ambiguous in the same sense that log a is. To obtain a value of ab, first choose a logarithm of a; call it log a. Such a choice may be the principal value Log a (the default, if no other specification is given), or perhaps a value given by some other branch of log z fixed in advance. Then, using the complex exponential function one defines

OpenStudy (perl):

a^b*

OpenStudy (perl):

so we have a^b = e^ ( ln (a^b)) = e^ ( b ln a )

OpenStudy (anonymous):

So in one word, we can say the definition of a^b=e^ ( b ln a )

OpenStudy (anonymous):

is that so?

OpenStudy (perl):

taking the principal value of ln a

OpenStudy (anonymous):

Ok, so the problem is solved is solved, I suppose, isn't it?

OpenStudy (perl):

and be careful when you read this, Log z is a special function

OpenStudy (perl):

it is not log base 10 , lol

OpenStudy (anonymous):

I know, its ln, right?

OpenStudy (perl):

err, im wrong , i dont really kno

OpenStudy (perl):

hey, do you have a pdf or website on complex math?

OpenStudy (perl):

log is base ten generally

OpenStudy (anonymous):

I am not getting you, what I meant to say, is the log to the base e is called ln

OpenStudy (perl):

oh yes

OpenStudy (perl):

where is your pdf on complex math!

OpenStudy (perl):

or your class has some online reading i bet

OpenStudy (perl):

quick before i fall asleep

OpenStudy (anonymous):

I don't have any pdf, I follow books, recommended to me by my profs

OpenStudy (anonymous):

Ok thank you so much for the help. Hope you will help me again in future. Is there a way by which I can direct you to my questions?